OpenStudy (anonymous):

write the equation of the line that contains the given points... A(6,-6) B(-7,7)

OpenStudy (anonymous):

are you sure you cant do this by yourself?

OpenStudy (anonymous):

do you know how to find the slope?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so, what is the slope?

OpenStudy (anonymous):

it's -13

OpenStudy (anonymous):

i would check that again

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

thats the one

OpenStudy (anonymous):

i emant to say

OpenStudy (anonymous):

i had -13/13, which is -1

OpenStudy (anonymous):

k hold on

OpenStudy (anonymous):

now, you have to use the point slope formula, using any one of those points as (x1,y,1) and plug that into : y-y1=m(x-x1)

OpenStudy (anonymous):

so use x1 or y1 and plug that in?

OpenStudy (anonymous):

choose one of your ordered pairs and plug that into the formula

OpenStudy (anonymous):

?

OpenStudy (anonymous):

for example: (6,-6)=(X1,Y1)

OpenStudy (anonymous):

so i wanna use x1 and x2

OpenStudy (anonymous):

y-(-6)=-1(x-6)

OpenStudy (anonymous):

6,-6=6,-7??????

OpenStudy (anonymous):

now just simplify that

OpenStudy (anonymous):

0=-1?

OpenStudy (anonymous):

y-(-6)=-1(x-6), once you simplify this, it will give you the equation of the line that contians both given points

OpenStudy (anonymous):

don't get it....

OpenStudy (anonymous):

y-y1=m(x-x1)----->this is the point slope equation of a line m---->slope (x1,y1)------->one of the pairs of points you were given

OpenStudy (anonymous):

so it would be y--7=m(x-6)???

OpenStudy (anonymous):

no, you only use one of your points

OpenStudy (anonymous):

example?

OpenStudy (anonymous):

in your case: m(slope)=-1 (x1,y1)=(6,-6) y-y1=m(x-x1) thus, plugging in our values: y-(-6)=-1(x-6) Now simplify that: y+6=-1x+6 y=-1x, that our line

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

you can hit the good answer button if you feel like thanking me for my help